OpenStudy (anonymous):

Given a polynomial f(x), if (x + 5) is a factor, what else must be true?

OpenStudy (anonymous):

@ganeshie8 @Abhisar @dan815 @Abb0t

OpenStudy (anonymous):

@peachpi

OpenStudy (danjs):

zeroes of the polynomial

OpenStudy (anonymous):

A.f(0) = 5 B. f(0) = −5 c. f(5) = 0 d. f(−5) = 0

OpenStudy (danjs):

If x-5 factors out of the thing, then you have f(x) = (x-5)(something left) if x-5=0 then....

OpenStudy (danjs):

Sorry it is f(x) = (x+5)(something else) Put in X= -5 f(-5) = 0*(Something) = 0

OpenStudy (anonymous):

Is it A?

OpenStudy (danjs):

If (x+5) is a factor, then if you let X=-5 , that factor will be zero, and multiply the rest of the function by zero f(-5) = ...

OpenStudy (anonymous):

oh lol it's D!!

OpenStudy (anonymous):

0

OpenStudy (danjs):

right, you see

OpenStudy (anonymous):

so it is D?

OpenStudy (danjs):

Right, Say the function is f(x) = (x+5)(x^3+6x-4) The x+5 is factored out of the thing... if that quantity is zero, (when x=-5 , f(x)=f(-5) then the function will also be zero at that location

OpenStudy (anonymous):

If x = 1 is a zero of the polynomial function f(x) = 3x3 − 8x2 + 3x + 2, which of the following is another zero of f(x)?

OpenStudy (danjs):

If x=1 is a zero, then you can probably say (x-1) will be a factor of f(x)

OpenStudy (anonymous):

So would it be -1?

OpenStudy (danjs):

no, that means you can say f(x) = (x-1)*(another factor)

OpenStudy (danjs):

When "another factor" = 0 , that will be you other zero(s)

OpenStudy (danjs):

The zeroes are where y=f(x) = 0, crosses x axis

OpenStudy (anonymous):

@DanJS so it is 2?